Given two vectors A = -2.00i + 4.00 j + 4.00 k and
Answer #1

\large \overset{}{\rightarrow}A=-2.0i+4.0j+4.0k

\large \overset{}{\rightarrow}B=4.0i+1.0j-3.0k

Part 1

\large I\overset{}{\rightarrow}AI=\sqrt{2^2+4^2+4^2}=6unit

Part 2

\large I\overset{}{\rightarrow}BI=\sqrt{4^2+1^2+3^2}=5.1unit

Part 3

\large \overset{}{\rightarrow}A-\overset{}{\rightarrow}B=(-2.0i+4.0j+4.0k)-(4.0i+1.0j-3.0k)

\large \overset{}{\rightarrow}A-\overset{}{\rightarrow}B=(-6.0i+3.0j+7.0k)

\large x=-6,y=3,z=7

Part D

IA-Bl = V62 + 32 + 72 = 9.695 unit

\large \overset{}{\rightarrow}B-\overset{}{\rightarrow}A=(4.0i+1.0j-3.0k)-(-2.0i+4.0j+4.0k)

B _ A = (6.01-3.0jー7.0k)

Part E

\large I\overset{}{\rightarrow}B-\overset{}{\rightarrow}AI=\sqrt{6^2+3^2+7^2}=9.695unit

Yes it is same